The Cheyenne Hotel in Big Sky, Montana, has accumulated records of the total electrical costs of the hotel and the number of occupancy-days over the last year. An occupancy-day represents a room rented for one day. The hotel’s business is highly seasonal, with peaks occurring during the ski season and in the summer. Month Occupancy-Days Electrical Costs January 2,710 $ 5,270 February 3,600 $ 6,205 March 860 $ 2,150 April 2,170 $ 4,350 May 4,200 $ 7,160 June 1,530 $ 3,825 July 4,110 $ 7,050 August 4,060 $ 6,925 September 1,980 $ 4,090 October 1,160 $ 2,900 November 1,210 $ 3,025 December 2,480 $ 4,910 Required: 1. Using the high-low method, estimate the fixed cost of electricity per month and the variable cost of electricity per occupancy-day. (Do not round your intermediate calculations. Round your Variable cost answer to 2 decimal places and Fixed cost element answer to nearest whole dollar amount.) 2. What other factors in addition to occupancy-days are likely to affect the variation in electrical costs from month to month? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answer and double click the box with the question mark to empty the box for a wrong answer. Any boxes left with a question mark will be automatically graded as incorrect.) Systematic factors like guests, switching off fans and lights. unchecked Number of days present in a month. unchecked Income taxes paid on hotel income. unanswered Seasonal factors like winter or summer. unanswered Fixed salary paid to hotel receptionist. unanswered
Answer #1
A)
Occupancy Electrical
Days costs
High activity level 4,200 7,160
low activity level 860 2,150
Change 3,340 5,010
Variable cost per unit 1.5
Fixed cost element 860
Working notes:
Variable cost per unit = 5010/3340 = 1.5 per unit
Fixed cost                      = 7160 - 4200*1.5= 860
B) Seasonal factors winter or summer
Systematic factors like guests,switching off fans and lights.